APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Saturday August 18, 2001 (MJD 52140) ---=== OBSERVING TEAM ===--- Swing: Pete Newman (APO) Night: Mike Harvanek (APO) Atsuko Nitta Kleinman (APO) ---=== OBSERVING PLAN ===--- Science. ---=== OBSERVING SUMMARY ===--- We finished one plate. The humidity readings we get from "display weather" showed some wild readings around 3am. We filed a PR on this since we rely heavily on the humidity readings -- especially this dark run since we are not quite out of the rainy season yet. ccd03's ln2temp readings show larger scatter than the others, but neither spectrograph stopped us from getting data tonight. ---=== OBSERVING LOG ===--- SWING SHIFT LOG =============== IOP servers were restarted under v3_53_6. All spectro CCDs were saturated on the first bias exposure (#10569), presumably from the problems reported last night. SoS did not report the first bias. The second bias exposure looked normal on the displays, but was SoS reported high 98th percentile values on sp1. Further biases were taken until the CCDs were cleared. During this process, grabInstruments failed yet again because port 30410 (the first UDP port allocated by the operating system?) was in use by an fsaoImage process that had been started with startMon. This failure mode seems to be occuring more often this season. The relevant PR (SOP/2217) was updated and its priority raised. At the end of the imager checkout bias drift, watcher reported a VME error on AD2: "Status Inactive; Run 2498; Frame 91; dead (Frame=91 Row=1361: Aug 18 17:22:23 MDT). A second short bias drift was taken, and the error was not seen. (What does DEAD mean?!) Note that frame 91 is not the last frame in that drift. NIGHT ====== With the current weather/humidity, when there is an LN2 fill, the dewars and the back of the spectrographs become very cold and get a lot of condensation which turns into ice . Sometime after the fill, the ice melts and we get water on the floor. Mike wiped off a lot of water today during the afternoon checkout. He also noted that almost every afternoon when we get the rain/thunders, the humidity inside the enclosure rise up to 90% or so. There was no condensation on the dog house/telescope, but the shroud we use to cover the laptop next to the scope is damp for the past few days from the high humidity weather. Once we were sure that the humidity was stable, we opened with cartridge 2. We went to field and found our guide stars! No fk5 was needed! Yupeee! As of last night and the night before, when we tried to do "preCalib", we got the 30410 udp socket in use error and could not take any exposures. We had to kill Mike's PGplot and reissue grabInst. Mike operated without the PGplot after this -- since all info is in the scrolling sop window, he used that to get scale and FWHM info. M2's axis E problem is still with us, but we just ignore and operate as Russell told us to do. TPM showed the 3rd ID switch as "disconnect". We have installed the cartridge with over 1000lb and ID switch should have registered even with less weight. TPM showed correct ID for cartridge 4 which was mounted last night and left it on until we changed it right before we opened tonight. As soon as we started science exposure, clouds rolled in. We hoped the weather to improve --- we ended up closing 23:30 since the humidity was starting to creep up again. After we closed, we noticed that the inst ID switch on TPM were all registering cart2 as it should. Something flaky with the switch? The weather condition improved and we opened at 00:45 with cartridge 4 on. Around 1:45 or so, we saw "ln2 empty"error on sp2. When we checked the ln2 temperature, ccd03 had ln2temp~79 and it showed up in watcher in red. The next reading was 89 and black. BTW, we noticed that Watcher's "active" button did not actively update the values and we had to force updates. PR was filed -- we like the "active" feature and want it back. The Watcher plot of ln2temp for ccd03 shows much larger scatter than other CCDs even though most of the time the values are within the nominal value. So the problem we saw last night might just be waiting to happen any minute...... Due to the climbing humidity, we closed after we finished cart 4. While we were waiting for the humidity to improve, we saw curious reports of (diff)=(temp)-(dew point)=-7.9C in the "weather display" on Galileo! The next reading went back to somewhat more reasonable number of 4 C. After this, we noticed that the display weather was reporting "Diff" that was anywhere from -11 to 6. We rely on these numbers reported by the display weather (and feeling the structure/doghouse for condensation) to judge if we should open/close. Considering the importance of the humidity/temperature/dew point readings to our operation, high critical PR was filed. We almost opened around 3:30, but we decided against it after we felt the railings/doors/buildings. We ran endNight with success. I had no problem pasting tonight using logTool. But curiously, I saw no "default" values in the logTool either ( e.g. software version numbers, plan etc.). I am on the same terminal that I was two nights ago when I had lots of frustrating pasting experience. ---=== IMAGING RUN SUMMARY ===--- IMAGING DATA DURING CHECKOUT ============================ Run Stripe Strip Flavor Lambda Last Frame # Begin End ------------------------------------------------------------------ 2498 100 O bias 14.70 27.22 98 doghouse 2499 100 O bias 31.75 34.66 33 doghouse ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- ---=== LTMATCH RESULTS ===--- ---=== SPECTROSCOPY DATA SUMMARY ===--- SPECTROSCOPY DATA DURING CHECKOUT ================================ UT Exp Time flavor comment =============================================== ----- sequence 10569, plate -9999 ------- 22:37 10569 0.0 bias {{}} Saturated 22:40 10570 0.0 bias {{}} Out of limits ----- sequence 10571, plate 375 ------- 23:19 10571 0.0 bias {{}} Within limits SPECTROSCOPY DATA AT NIGHT ============================== logTool sometimes report the same exposure multiple times as one can see here. ----- sequence 10573, plate 634 ------- 04:05 10573 10.0 flat pre-flat ----- sequence 10573, plate 634 ------- 04:05 10573 10.0 flat pre-flat ----- sequence 10573, plate 634 ------- 04:05 10573 10.0 flat pre-flat ----- sequence 10573, plate 634 ------- 04:05 10573 10.0 flat pre-flat 04:07 10574 2.0 arc pre-arc 04:07 10574 2.0 arc pre-arc 04:07 10574 2.0 arc pre-arc 04:07 10574 2.0 arc pre-arc 04:26 10575 900.0 target science 04:26 10575 900.0 target science 04:26 10575 900.0 target science 04:26 10575 900.0 target science 04:44 10576 900.0 target science 04:44 10576 900.0 target science 04:44 10576 900.0 target science 04:44 10576 900.0 target science 05:02 10577 901.0 target science 05:02 10577 901.0 target science 05:02 10577 901.0 target science 05:02 10577 901.0 target science ----- sequence 10578, plate 375 ------- 07:15 10578 10.0 flat pre-flat ----- sequence 10578, plate 375 ------- 07:15 10578 10.0 flat pre-flat ----- sequence 10578, plate 375 ------- 07:15 10578 10.0 flat pre-flat 07:17 10579 2.0 arc pre-arc 07:17 10579 2.0 arc pre-arc 07:17 10579 2.0 arc pre-arc 07:17 10579 2.0 arc pre-arc 07:41 10580 901.0 target science 07:41 10580 901.0 target science 07:41 10580 901.0 target science 07:58 10581 900.0 target science 07:58 10581 900.0 target science 07:58 10581 900.0 target science 08:16 10582 900.0 target science 08:16 10582 900.0 target science 08:16 10582 900.0 target science 08:23 10583 240.0 smear {{Smear Pattern}} 08:23 10583 240.0 smear {{Smear Pattern}} 08:23 10583 240.0 smear {{Smear Pattern}} 08:23 10583 240.0 smear {{Smear Pattern}} ---=== TELESCOPE OFFSETS AND SCALE I ===--- 22:05 4 375 998193933 0.002538 0.000606 0.010310 0.999950 01:23 4 375 998205806 0.004572 0.000844 0.004901 1.000000 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- FNAL: JL4291 APO: 4292 ---=== FOCUS LOG ===--- 22:06 4 375 0.99995 493 1511 -100 38 43.9 14.0 5 90 1.2" 01:23 4 375 1.00000 0 1010 -150 5 57.1 14.2 0 247 1.4" ---=== WEATHER LOG ===--- 22:00 57 48 6 80 (E) 1436 - 22:32 57 48 2 84 (E) 1588 - 23:03 57 48 5 106 (ESE) 1448 - 23:34 56 48 10 123 (ESE) 1383 - 00:07 56 48 5 96 (E) 1337 - 01:10 58 42 0 84 (E) 884 - 01:43 56 45 0 226 (SW) 1165 - 02:15 56 45 1 268 (W) 1347 - 02:47 57 44 0 33 (NNE) 1386 - 03:19 55 47 0 302 (WNW) 1566 - 03:51 55 47 1 161 (SSE) 1396 - ---=== TELESCOPE STATUS ===--- Doors open: 20:00, 00:45 Doors open: 21:30, 02:30 Fans on : 20:00, 00:45 Fans off : 21:30, 02:30 Enclosure open: 21:30, 00:45 Enclosure close: 23:30, 02:30 telescope stowed at (121,30) with cartridge 3 on. ---=== SOFTWARE USED ===--- iop/sop v3_53_6 tccmon v1_6 sdssProcedure v1_22 Watcher v2_9 TPM v2_4_1 MCP v5_3_1 astroda v14_36 TCC 2.5.5 ---=== MIRROR NUMBERS ===--- im> Time dpTemp dewPt Press TmpOut HumOut Windd Winds 19:39 58.3 50.7 21.68 75 184.0 4.6 14 PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A: 0.0720 0.0710 Axial B: -0.0880 -0.0980 Axial C: 0.0670 0.0610 Trans D: 4.1890 4.1800 Lateral E: 1.9330 1.9365 Lateral F: 0.4335 0.4370 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES primDesOrient: 0.00 -11.80 23.00 1256.90 642.10 primOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.0 Air temp: 14.6 C Altitude: 30.44 MIGS TONIGHT NOMINAL Axial A: -1.2940 -1.2900 Axial B: -1.3020 -1.3150 Axial C: -1.3120 -1.3060 Trans D: 0.1360 0.1890 <-- at the edge of acceptable value GALILS Commanded: 1583850. 1583850. 1583850. 2600. 3800. Actual: 1583926. 1583857. 1583851. 2600. 3800. SETMIR VALUES secDesOrient: 1257.00 0.00 0.00 0.00 -175.00 secOrient: 1257.03 0.01 -0.02 -0.02 -175.13 ---=== PROBLEMS IN DETAIL ===---